# From Distance to Angle: One-Shot Detection Under Additive White Cauchy Noise

### 10 Apr 2026 v1

@@ Abstract
We study one-shot detection under additive white Cauchy noise (AWCN) using finite constellations, with emphasis on the geometric mechanisms governing symbol-level reliability. Under isotropic Cauchy noise, the maximum-likelihood rule induces the same Euclidean Voronoi decision regions as in the Gaussian case, so the distinction lies not in the decision geometry itself but in how probability mass is distributed over these fixed regions. In the small-noise regime, we derive a reciprocal distance-spectrum upper bound for the symbol error probability, showing that reliability retains a longer-range dependence on the global constellation geometry than under additive white Gaussian noise. In the large-noise regime, we prove that the correct-decision probability converges to a limit determined solely by the angular measure of the associated Voronoi recession cone. These results formalize a regime-dependent transition from distance-based to angle-based reliability descriptors under heavy-tailed noise. The theory is further illustrated through an asymmetric four-point example exhibiting geometric collapse and a standard 4QAM sanity check.
@@ I Introduction
Additive white Gaussian noise (AWGN) has long served as the canonical model for detection and constellation design, under which reliability is largely governed by minimum-distance geometry. This viewpoint underlies much of modern digital communication theory [9], particularly in the small-noise regime. Recent studies, however, have identified communication and sensing scenarios in which noise exhibits heavy-tailed, non-Gaussian behavior, rendering Gaussian-based design principles suboptimal [10, 6]. Such heavy-tailed noise models have also attracted recent interest in information-theoretic studies [1, 15, 8].
Specifically, the Cauchy model has emerged as a prototypical member of the $\alpha$-stable family [6], offering a mathematically rigorous yet tractable departure from the Gaussian assumption. While its infinite variance and polynomial tail decay complicate traditional Law-of-Large-Numbers-based analysis [15], the model is far from a mere theoretical abstraction. It arises naturally in practical systems, including massive MIMO [4] and diffusion-based molecular communication [2, 5], where nearest-neighbor geometry no longer suffices to characterize reliability. This necessitates identifying the alternative geometric descriptors under heavy-tailed noise, an area partially explored in [11, 13, 12].
On the other hand, recent findings [8] reveal that channels subject to AWCN can sustain positive information rates under a second-moment input (energy) constraint. By positioning AWCN as a meaningful benchmark directly comparable to AWGN under identical energy normalization, these results motivate a timely re-examination of finite-constellation detection and design [14, 9] in the presence of Cauchy noise.
To address this need, this paper formally develops a geometric asymptotic theory for one-shot detection over AWCN channels. Our focus is not on modifying the decision rule itself, which remains Euclidean under isotropic Cauchy noise, but on identifying the geometric descriptors that govern reliability across different noise regimes. In the small-noise regime, we derive a reciprocal distance-spectrum upper bound showing that one-shot reliability retains a longer-range sensitivity to the global constellation geometry than in the Gaussian case. In the large-noise regime, we prove that bounded distance information disappears asymptotically and that the correct-decision probability converges to a limit determined solely by the angular measure of the associated Voronoi recession cone (see [7]). Taken together, these results provide a theorem-based account of how one-shot reliability shifts from distance-dominated behavior to angle-dominated behavior under heavy-tailed noise.
@@
The remainder of this paper is organized as follows. Section II introduces the system model and problem statement. Section III presents the small-noise distance-spectrum bound. Section IV establishes the large-noise angular convergence theorem and the associated geometric-collapse criterion. The main results are further illustrated through an asymmetric four-point worked example in Appendix A and a standard 44QAM sanity check in Appendix B. Conclusions are drawn in Section V.
@@ II-A Vector Cauchy Channels
where $\bm{x}\in\mathcal{C}\subset\mathbb{R}^{d}$ is the transmitted signal vector from a finite constellation subject to an average energy constraint $\mathbb{E}[\|\bm{x}\|^{2}]\leq P$. The additive noise $\bm{n}\in\mathbb{R}^{d}$ is modeled as an isotropic $d$-dimensional Cauchy random vector with probability density function (PDF) (see [10, 15]),
@@
where $\gamma>0$ is the scale parameter controlling the distribution’s width. Unlike Gaussian noise, the Cauchy distribution lacks finite moments; thus, standard signal-to-noise ratio (SNR) is technically undefined. We instead use the normalized proxy $P/\gamma^{2}$ to characterize system performance (see [8]).
Under this spherically symmetric model, the ML decoding rule minimizes the negative log-likelihood, which simplifies gracefully to the standard Euclidean distance metric. Namely,
@@
Because the isotropic Cauchy density is monotonically decreasing with the Euclidean norm, the exact decision boundaries induced by (3) correspond to Euclidean Voronoi cells (see [7]), entirely independent of $\gamma$. Under equiprobable signaling, the average symbol error probability is formally given by
$P_{e}=1-\frac{1}{M}\sum_{\bm{x}\in\mathcal{C}}\int_{V(\bm{x})}f_{\bm{N}}(\bm{y}-\bm{x})\mathrm{d}\bm{y}.$ (4)
@@ II-B The Combining Gain Paradox
In traditional AWGN vector channels, linear aggregation strategies (e.g., maximum ratio combining) yield a genuine averaging effect, so reliability typically improves with the number of spatial dimensions or channel paths [9]. Under AWCN, this self-averaging mechanism breaks down: the Cauchy law has neither finite mean nor finite variance, so the usual Law-of-Large-Numbers (LLN) intuition no longer applies. More fundamentally, because the Cauchy distribution is $1$-stable (see [10, 6]), any linear combination of independent Cauchy variables remains Cauchy. Consequently, linear combining does not concentrate the effective noise distribution, but merely produces another scalar Cauchy projection, thereby losing the classical multidimensional combining gain.
Because the exact $d$-dimensional Cauchy integral in (4) lacks a closed-form solution over arbitrary Voronoi polytopes, resolving this paradox requires moving beyond linear operations and standard distance formulations. The specific objective of this paper is to geometrically reinterpret the error integral (4). By establishing how symbol reliability transitions from distance-based to angle-based mechanisms, we aim to identify the geometric descriptors that govern one-shot detection under AWCN and that may inform future geometry-aware receiver design.
@@ III-B Main Result: Small-Noise Bound and Asymptotics
$=\int_{d_{ij}/2}^{\infty}\frac{1}{\pi\gamma}\frac{1}{1+(z/\gamma)^{2}}\,\mathrm{d}z.$
@@
Theorem 1 clarifies an important contrast between Gaussian and Cauchy detection. Under AWGN, pairwise error terms decay exponentially with distance, so the nearest-neighbor structure typically dominates the small-noise behavior. Under AWCN, however, the pairwise comparison term in (9) satisfies
@@ IV-B Main Result: Large-Noise Convergence
$P_{c}(x;\gamma)=\int_{V(x)-x}c_{d}\,\gamma^{-d}\left(1+\frac{\|n\|^{2}}{\gamma^{2}}\right)^{-\frac{d+1}{2}}\,\mathrm{d}n.$ (47)
@@
$P_{c}(x;\gamma)=\int_{(V(x)-x)/\gamma}c_{d}\,(1+\|u\|^{2})^{-\frac{d+1}{2}}\,\mathrm{d}u.$ (48)
@@
$\lim_{\gamma\to\infty}P_{c}(x;\gamma)=\int_{\mathsf{K}(x)}c_{d}(1+\|u\|^{2})^{-\frac{d+1}{2}}\,\mathrm{d}u.$ (55)
@@
$1=\int_{\mathbb{R}^{d}}c_{d}(1+\|u\|^{2})^{-\frac{d+1}{2}}\,\mathrm{d}u=\sigma_{d-1}(\mathbb{S}^{d-1})\,I_{d}.$ (58)
@@
V Conclusion

### 20 Apr 2026 v2

# From Distance to Angle: One-Shot Detection Under Isotropic Multivariate Cauchy Noise

@@ Abstract
We study one-shot detection under isotropic multivariate Cauchy noise using finite constellations, with emphasis on the geometric mechanisms governing symbol-level reliability. Under isotropic Cauchy noise, the maximum-likelihood rule induces the same Euclidean Voronoi decision regions as in the Gaussian case, so the distinction lies not in the decision geometry itself but in how probability mass is distributed over these fixed regions. In the small-noise regime, we derive a reciprocal distance-spectrum upper bound for the symbol error probability, showing that reliability retains a longer-range dependence on the global constellation geometry than under additive white Gaussian noise. In the large-noise regime, we prove that the correct-decision probability converges to a limit determined solely by the angular measure of the associated Voronoi recession cone. These results formalize a regime-dependent transition from distance-based to angle-based reliability descriptors under heavy-tailed noise. The theory is further illustrated through an asymmetric four-point example exhibiting geometric collapse, a standard 4QAM sanity check, and finite-$\gamma$ numerical validation for both asymptotic regimes.
@@ I Introduction
Additive white Gaussian noise (AWGN) has long served as the canonical model for detection and constellation design, under which reliability is largely governed by minimum-distance geometry. This viewpoint underlies much of modern digital communication theory [13], particularly in the small-noise regime. Recent studies, however, have identified communication and sensing scenarios in which noise exhibits heavy-tailed, non-Gaussian behavior, rendering Gaussian-based design principles suboptimal [14, 10]. Such heavy-tailed noise models have also attracted sustained interest across information theory, signal processing, and communication systems [2, 16, 12].
Specifically, the Cauchy model has emerged as a prototypical member of the $\alpha$-stable family [10], offering a mathematically rigorous yet tractable departure from the Gaussian assumption. While its infinite variance and polynomial tail decay complicate traditional concentration-based analysis (e.g., via the Law-of-Large-Numbers (LLN)) [16], the model is far from a mere theoretical abstraction. It arises naturally in practical systems, including massive MIMO [7] and diffusion-based molecular communication [5, 8]. At a broader level, existing work on stable and Cauchy noise has progressed along several largely separate directions. In information theory, the adopted input constraint often changes with the characteristic exponent: the Gaussian case is classically studied under a second-moment constraint, the Cauchy case has been analyzed under logarithmic cost [3], and more general $\alpha$-stable channels have also been treated under alternative moment constraints to obtain tractable capacity bounds [1]. In signal processing, impulsive environments have motivated generalized information and estimation tools beyond the Gaussian second-order framework [4]. In communication systems, heavy-tailed laws also arise from concrete channel models rather than from abstract distributional choices alone; in particular, Cauchy-type behavior can naturally emerge in molecular communication settings such as first-arrival-position channels [8, 9].
These developments point to a common missing piece. Although recent findings [12] show that memoryless scalar Cauchy-noise channels can sustain positive information rates under a second-moment input (energy) constraint, and thus support a renewed power-constrained viewpoint in the Cauchy regime, a multivariate geometric theory for finite-constellation detection under isotropic multivariate Cauchy noise remains undeveloped. In particular, from the perspective of symbol-level detection, it is still unclear which geometric descriptors replace the classical minimum-distance paradigm once the noise law becomes heavy-tailed. This motivates a timely re-examination of finite-constellation detection and design (cf. [15, 13]) in the presence of Cauchy noise.
To address this need, this paper develops a geometric asymptotic theory for one-shot detection over channels corrupted by isotropic multivariate Cauchy noise. Our main contributions are summarized as follows:
  • We formulate one-shot finite-constellation detection under isotropic multivariate Cauchy noise and clarify that, although the maximum-likelihood rule induces the same Euclidean Voronoi decision regions as in the Gaussian case, the associated reliability laws are governed by fundamentally different geometric mechanisms.
  • In the small-noise regime, we derive a reciprocal distance-spectrum upper bound showing that one-shot reliability retains a longer-range sensitivity to the global constellation geometry than in the Gaussian setting.
  • In the large-noise regime, we prove that bounded distance information disappears asymptotically and that the correct-decision probability converges to a limit determined solely by the angular measure of the associated Voronoi recession cone (cf. [11]), thereby identifying an angle-based reliability descriptor and a corresponding geometric-collapse criterion.
  • We further support this regime-dependent transition through explicit worked examples and finite-$\gamma$ numerical validation, showing how the theory manifests in both asymmetric (see Appendix A) and symmetric (see Appendix B) constellations.
@@
The remainder of this paper is organized as follows. Section II introduces the system model and problem statement. Section III presents the small-noise distance-spectrum bound. Section IV establishes the large-noise angular convergence theorem and the associated geometric-collapse criterion. Section V provides finite-$\gamma$ numerical validation for both asymptotic regimes. The main results are further illustrated through an asymmetric four-point worked example in Appendix A and a standard 44QAM sanity check in Appendix B. Conclusions are drawn in Section VI.
@@ II-A Vector Cauchy Channels
where $\bm{x}\in\mathcal{C}\subset\mathbb{R}^{d}$ is the transmitted signal vector from a finite constellation subject to an average energy constraint $\mathbb{E}[\|\bm{x}\|^{2}]\leq P$. The additive noise $\bm{n}\in\mathbb{R}^{d}$ is modeled as an isotropic $d$-dimensional Cauchy random vector with probability density function (PDF) (cf. [14, 16]),
@@
where $\gamma>0$ is the scale parameter controlling the distribution’s width. Here, the noise is modeled as a single isotropic multivariate Cauchy vector in $\mathbb{R}^{d}$, rather than as independent componentwise Cauchy coordinates. Unlike Gaussian noise, the Cauchy distribution lacks finite moments; thus, standard signal-to-noise ratio (SNR) is technically undefined. We instead use the normalized proxy $P/\gamma^{2}$ to characterize system performance (cf. [12]).
In the Gaussian setting, isotropy and componentwise i.i.d. structure lead to the same multivariate law. In the Cauchy setting, however, these two notions are no longer equivalent: the isotropic multivariate Cauchy law is radially symmetric, whereas independent Cauchy coordinates define a different joint distribution. This distinction matters here because our analysis is carried out under an isotropic multivariate Cauchy observation model rather than a componentwise independent one. It also places the present work in contrast with the memoryless scalar Cauchy-noise setting studied in [12].
Under this spherically symmetric model, the ML decoding rule minimizes the negative log-likelihood, which reduces to the standard Euclidean distance metric. Namely,
@@
Because the isotropic Cauchy density is monotonically decreasing with the Euclidean norm, the exact decision boundaries induced by (3) correspond to Euclidean Voronoi cells (cf. [11]), entirely independent of $\gamma$. Under equiprobable signaling, the average symbol error probability is formally given by
$P_{e}=1-\frac{1}{M}\sum_{\bm{x}\in\mathcal{C}}\int_{V(\bm{x})}f_{\bm{N}}(\bm{y}-\bm{x})~\mathrm{d}\bm{y}.$ (4)
@@ II-B The Combining Gain Paradox
In traditional AWGN vector channels, linear aggregation strategies (e.g., maximum ratio combining) often yield an averaging gain under standard finite-variance assumptions, so reliability can improve with the number of spatial dimensions or channel paths [13]. Under isotropic multivariate Cauchy noise, this self-averaging intuition breaks down: the Cauchy law has neither finite mean nor finite variance, so the usual LLN viewpoint no longer applies. More fundamentally, under the isotropic multivariate Cauchy model, every one-dimensional projection onto a unit direction remains a scalar Cauchy random variable with the same scale parameter (cf. [14, 10]). Consequently, linear combining does not reduce the effective noise scale in the usual Gaussian sense, but instead produces another scalar Cauchy projection, thereby removing the classical concentration effect associated with multidimensional combining.
Because the exact $d$-dimensional Cauchy integral in (4) lacks a closed-form solution over arbitrary Voronoi polytopes, understanding this contrast calls for a viewpoint beyond linear averaging arguments alone. The specific objective of this paper is therefore to geometrically reinterpret the error integral (4). By showing how symbol reliability transitions from distance-based to angle-based behavior, we aim to identify the geometric descriptors that govern one-shot detection under the isotropic Cauchy model and that may inform future geometry-aware receiver design.
@@ III-B Main Result: Small-Noise Bound and Asymptotics
$=\int_{d_{ij}/2}^{\infty}\frac{1}{\pi\gamma}\frac{1}{1+(z/\gamma)^{2}}~\mathrm{d}z.$
@@
Theorem 1 clarifies an important contrast between Gaussian and Cauchy detection. Under AWGN, pairwise error terms decay exponentially with distance, so the nearest-neighbor structure typically dominates the small-noise behavior. Under the isotropic Cauchy model, however, the pairwise comparison term in (9) satisfies
@@ IV-B Main Result: Large-Noise Convergence
$P_{c}(x;\gamma)=\int_{V(x)-x}c_{d}\,\gamma^{-d}\left(1+\frac{\|n\|^{2}}{\gamma^{2}}\right)^{-\frac{d+1}{2}}~\mathrm{d}n.$ (47)
@@
$P_{c}(x;\gamma)=\int_{(V(x)-x)/\gamma}c_{d}\,(1+\|u\|^{2})^{-\frac{d+1}{2}}~\mathrm{d}u.$ (48)
@@
$\lim_{\gamma\to\infty}P_{c}(x;\gamma)=\int_{\mathsf{K}(x)}c_{d}(1+\|u\|^{2})^{-\frac{d+1}{2}}~\mathrm{d}u.$ (55)
@@
$1=\int_{\mathbb{R}^{d}}c_{d}(1+\|u\|^{2})^{-\frac{d+1}{2}}~\mathrm{d}u=\sigma_{d-1}(\mathbb{S}^{d-1})\,I_{d}.$ (58)
@@
(a) Average symbol error probability in the small-$\gamma$ regime for the asymmetric four-point constellation in Appendix A. The Monte Carlo estimate is compared with the exact upper bound from Theorem 1 and with its leading-order asymptotic approximation. Over the displayed range in (71), the Monte Carlo curve remains below the exact bound, while the exact bound and the leading-order asymptotic are nearly indistinguishable. This shows that the reciprocal-distance-spectrum scaling predicted by Theorem 1 becomes visible already at finite small noise scales.
(b) Conditional symbol error probabilities in the same small-$\gamma$ regime for the asymmetric four-point constellation in Appendix A. For each transmitted symbol, the Monte Carlo estimates for $P_{1},P_{2},P_{3},$ and $P_{4}$ are shown together with the corresponding exact conditional bound and leading-order asymptotic line from Theorem 1. The ordering predicted in Appendix A from the reciprocal-distance sums is clearly observed: $P_{1}$ is the least reliable symbol, followed by $P_{2}$, whereas $P_{3}$ and $P_{4}$ are more reliable and nearly coincide. Because all four points have the same minimum distance, this figure highlights that under isotropic Cauchy noise, symbol reliability depends on the full reciprocal-distance geometry rather than on $d_{\min}$ alone.
Figure 1: Finite-$\gamma$ validation of the small-noise theory in Theorem 1 for the asymmetric four-point constellation from Appendix A. Panel (a) verifies that the average Monte Carlo symbol error probability remains below the exact bound in Theorem 1 and that the leading-order asymptotic already tracks the bound closely in the finite small-noise regime. Panel (b) shows that the same phenomenon persists at the symbol level and that the conditional reliability ordering is governed by the reciprocal-distance coefficients derived in Theorem 1.
@@
V Numerical Results
Unless otherwise stated, all Monte Carlo (MC) experiments in this section use equiprobable signaling, a fixed random seed rng(2026,’twister’), and $N_{\mathrm{mc}}=5\times 10^{5}$ trials for each value of $\gamma$, processed in batches of size $10^{5}$. The additive noise is generated as an isotropic multivariate Cauchy vector in $\mathbb{R}^{2}$ via
$N=\gamma\,\frac{G}{H},$ (69)
where $G\sim\mathcal{N}(0,I_{2})$ and $H\sim\mathcal{N}(0,1)$ are independent. In all simulations, maximum-likelihood decoding is implemented by the equivalent Euclidean nearest-neighbor rule in (3).
V-A Finite-$\gamma$ Validation of Theorem 1
We numerically validate Theorem 1 in the finite small-noise regime using the asymmetric four-point constellation from Appendix A,
$\mathcal{X}=\{(0,0),(1,0),(0,1),(0,-1)\}.$ (70)
For each
[…]
we estimate the symbol error probability by Monte Carlo simulation and compare it with both the exact union bound from Theorem 1 and its leading-order small-$\gamma$ asymptotic expression.
(a) Average correct-decision probability in the large-$\gamma$ regime for the asymmetric four-point constellation from Appendix A. The Monte Carlo estimate is compared with the angular limit predicted by Theorem 2, namely (76), and shown in the legend as “Angular limit (Thm. 2).” The simulated curve decreases steadily over the tested range and approaches the predicted plateau as the noise scale becomes large.
(b) Conditional correct-decision probabilities in the same large-$\gamma$ regime. The Monte Carlo estimates for $P_{1},P_{2},P_{3},$ and $P_{4}$ are shown explicitly in the legend, together with the corresponding angular (conditional) limits from Theorem 2 and Appendix A. The marker encoding is: black circles for $P_{1}=(0,0)$, blue squares for $P_{2}=(1,0)$, red upward triangles for $P_{3}=(0,1)$, and green downward triangles for $P_{4}=(0,-1)$. The numerical curves clearly approach the predicted limits in (75). In particular, the collapse of $P_{c}(P_{1};\gamma)$ toward zero provides direct numerical evidence for geometric collapse.
Figure 2: Finite-$\gamma$ validation of the large-noise theory in Theorem 2 for the asymmetric four-point constellation from Appendix A. Panel (a) shows that the average correct-decision probability approaches the angular limit $1/4$. Panel (b) shows that the corresponding symbol-level probabilities converge to their individual angular limits determined by the Voronoi recession cones. These results numerically confirm that, in the large-noise regime, reliability is governed by angular geometry rather than by bounded Euclidean distance information.
Fig. 1 summarizes the results. In Fig. 1(a), the Monte Carlo average symbol error probability remains strictly below the exact upper bound throughout the tested range, as expected. Moreover, the exact bound and its leading-order asymptotic approximation are nearly indistinguishable over the entire displayed small-$\gamma$ regime, indicating that the reciprocal-distance-spectrum law predicted by Theorem 1 emerges already at finite and practically observable noise levels. Although the upper bound is not numerically tight, it correctly captures the small-noise scaling and the relevant geometric dependence.
Fig. 1(b) further resolves the symbol-level behavior by plotting the conditional symbol error probabilities together with the corresponding exact conditional bounds and leading-order asymptotic lines. The ordering predicted in Appendix A is clearly observed:
$P_{e}(P_{1};\gamma)>P_{e}(P_{2};\gamma)>P_{e}(P_{3};\gamma)=P_{e}(P_{4};\gamma).$ (72)
Throughout the tested range, this numerical separation cannot be attributed to nearest-neighbor geometry alone, since all four points share the same minimum distance. Instead, it is consistent with the reciprocal-distance coefficients in Theorem 1, thereby illustrating that under isotropic Cauchy noise, even the high-reliability regime retains sensitivity to the global constellation geometry rather than only to $d_{\min}$.
Taken together, these simulations support two conclusions. First, Theorem 1 provides a finite-$\gamma$ upper bound with clear geometric content. Second, its leading-order reciprocal-distance descriptor already captures the correct qualitative reliability structure at the symbol level in the small-noise regime.
V-B Finite-$\gamma$ Validation of Theorem 2
We next validate the large-noise prediction in Theorem 2 using the same asymmetric four-point constellation from Appendix A,
$\mathcal{X}=\{(0,0),(1,0),(0,1),(0,-1)\}.$ (73)
For each
$\gamma\in\{$ 0.2, 0.3, 0.5, 0.7, 1, 1.5, $2,$03,05,07,1,15,2, (74)
$3,5,8,12,20,30\},$
we estimate the correct-decision probabilities by Monte Carlo simulation and compare them with the angular limits predicted by Theorem 2.
For this constellation, Appendix A gives the conditional large-noise limits
$\lim_{\gamma\to\infty}P_{c}(P_{1};\gamma)$ $=0,$ (75)
$\lim_{\gamma\to\infty}P_{c}(P_{2};\gamma)$ $=\frac{1}{4},$
$\lim_{\gamma\to\infty}P_{c}(P_{3};\gamma)=\lim_{\gamma\to\infty}P_{c}(P_{4};\gamma)$ $=\frac{3}{8},$
and hence the average correct-decision probability satisfies
$\lim_{\gamma\to\infty}P_{c}(\gamma)=\frac{1}{4}.$ (76)
Fig. 2 summarizes the numerical behavior. In Fig. 2(a), the Monte Carlo average correct-decision probability decreases steadily as $\gamma$ grows and approaches the predicted angular limit $1/4$. Although the convergence is naturally gradual, the trend is already clear on the finite-$\gamma$ range shown here, and the Monte Carlo values stabilize near the theorem-predicted plateau for the largest noise scales.
Fig. 2(b) further resolves the symbol-level behavior. The four conditional correct-decision probabilities approach distinct angular limits, exactly as predicted by Theorem 2 and Appendix A. In particular, the symbol $P_{1}=(0,0)$ becomes asymptotically unreliable, with its correct-decision probability tending to zero. This is the geometric-collapse phenomenon identified in Corollary 1. By contrast, $P_{2}$ approaches the nonzero limit $1/4$, while $P_{3}$ and $P_{4}$ approach the common limit $3/8$ and remain nearly indistinguishable throughout the simulation, consistent with the symmetry of the constellation.
These results provide finite-$\gamma$ numerical evidence for the large-noise geometric mechanism developed in Theorem 2. As the noise scale increases, bounded Euclidean offsets among constellation points lose influence, and the symbol-level reliability is governed instead by the angular measures of the corresponding Voronoi recession cones. Together with the small-noise validation in Section V-A, this experiment numerically supports the regime-dependent transition from distance-based to angle-based reliability descriptors under isotropic multivariate Cauchy noise.
VI Conclusion